OpenStudy (anonymous):

lim x->5 (4f(x))=7 is the lim x->5 (f(x))=7/4 ?

OpenStudy (anonymous):

@DebbieG help me please ? :)

OpenStudy (anonymous):

@pgpilot326 help me please ? :)

OpenStudy (jdoe0001):

I'd think so, yes \(\bf lim_{x\rightarrow 5} 4f(x) = 7 \implies lim_{x\rightarrow 5}4 = \cfrac{lim_{x\rightarrow 5} 7}{lim_{x\rightarrow 5} f(x)}\\ \implies lim_{x\rightarrow 5} f(x) = \cfrac{lim_{x\rightarrow 5} 7 }{lim_{x\rightarrow 5} 4}\)

OpenStudy (anonymous):

\[\lim_{x \rightarrow 5}4\cdot f(x) = 4\cdot \lim_{x \rightarrow 5}f(x)=7 \Rightarrow \lim_{x \rightarrow 5}f(x)= \frac {7}{4}\]